State Financial Corp. has three service departments (Administration, Communications, and Facilities), and two production departments (Deposits and Loans). A summary of costs and other data for each department prior to allocation of service department costs for the year ended December 31 follows.

Administration Communications Facilities Deposits Loans
Direct costs $ 240,000 $ 290,000 $ 260,000 $ 8,360,000 $ 4,700,000
Employee hours 25,000 33,000 19,000 445,000 343,000
Number of employees 9 18 5 230 150
Square footage occupied 4,600 12,700 5,500 243,300 204,600

  

The costs of the service departments are allocated on the following bases: Administration, employee-hours; Communications, number of employees; and Facilities, square footage occupied.

Required:

a. Assume that the bank elects to distribute service department costs to production departments using the direct method. What amount of Communications Department costs is allocated to the Deposits Department?

b. Assume the same method of allocation as in requirement (a). What amount of Administration Department costs is allocated to the Loans Department?

c. Assuming that the bank elects to distribute service department costs to other departments using the step method (starting with Facilities and then Communications), what amount of Facilities Department costs is allocated to the Communications Department?

d. Assume the same method of allocation as in requirement (c). What amount of Communication Department costs is allocated to Facilities?
Answer #1

a. Communications Department costs allocated to Deposits Department = $290000 x 230/(230 + 150) = $290000 x 230/380 = $175526.32

b. Administration Department costs allocated to Loans Department = $240000 x 343000/(445000 + 343000) = $240000 x 343000/788000 = $104467

c. Facilities Department costs allocated to Communications Department = $260000 x 12700/(4600 + 12700 + 243300 + 204600) = $260000 x 12700/465200 = $7098.02

d. None. Since the Facilities Department costs are allocated first, there is no back allocation of the Communication Department costs to the Facilities Department.

Kindly round off all answers as required since no instructions have been provided with the question regarding rounding off.